# MMD-Flagger: Leveraging Maximum Mean Discrepancy to Detect Hallucinations

### v1

@@ 2.2 Maximum Mean Discrepancy (MMD)
Maximum Mean Discrepancy (MMD) is a measure of distance between two probability distributions. It relies on a positive semi-definite kernel (Hofmann-2008, ; Smola-2009, ) to map the distributions into a new space, a Reproducing Kernel Hilbert Space (RKHS), where probability distributions are represented as data points (Muandet-2017, ). A large MMD value indicates significant differences between two distributions. We depict an intuitive presentation of the MMD in Figure 2.
@@
$\begin{split}{\rm MMD}_{k}^{2}(P,Q):=\ \mathbb{E}_{A,A^{\prime}\sim P}[k(A,A^{\prime})]+\mathbb{E}_{B,B^{\prime}\sim Q}[k(B,B^{\prime})]-2\mathbb{E}_{A\sim P,\ B\sim Q}[k(A,B)]\,,\end{split}$ (2)
@@
Empirical estimate. Given i.i.d. samples ${\bf A}=\{A^{1},\dots,A^{n}\}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}P$ and ${\bf B}=\{B^{1},\dots,B^{m}\}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Q$, one can estimate the MMD by replacing the expectations in Eq. (2) by the corresponding empirical averages:
$\begin{split}&\widehat{\rm MMD}^{2}_{U}({\bf A},{\bf B}):=\frac{1}{n(n-1)}\sum_{1\leq i\neq i^{\prime}\leq n}k(A^{i},A^{i^{\prime}})\\ &+\frac{1}{m(m-1)}\sum_{1\leq j\neq j^{\prime}\leq m}k(B^{j},B^{j^{\prime}})-\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}k(A^{i},B^{j}).\end{split}$ (3)
This is an unbiased estimate of the MMD Eq. (2), and converges at the rate $\mathcal{O}_{\mathbf{P}}(\min(n,m)^{-1/2})$ as $n,m\to\infty$ provided that $\sup_{x\in\mathbb{R}^{d}}k(x,x)<\infty$ (Gretton-2012a, , Theorem 10).
MMD in our setting. Recall that our goal is to measure the similarity between the outputs of $\text{LLM}_{\rm beam}$ and those of $\text{LLM}_{\rm sto}^{\tau}$. To do this, we consider $\text{LLM}_{\rm beam}$ and $\text{LLM}_{\rm sto}^{\tau}$ as two probability distributions, and use Eq. (3) to get an estimate of the MMD. Note that, since $\text{LLM}_{\rm beam}$ deterministically generates a sequence, it is simply a constant random variable.
Vector representation of outputs.
Since most kernels (especially the ones used in this work) take vector values as input, we need to convert outputs into vector representation. The most natural way is to resort to the word embedding matrix $E\in\mathbb{R}^{V\times D}$ used by our encoder-decoder. Thus each token sequence $y$ is associated with a sequence of embedding vectors
@@
We introduce two aggregations of $T$-length vector representations and convert them into a fixed-length vector $h$; avg and concat. The avg aggregation is computing the average over $T$ vectors, namely
@@
$h=[e_{1};e_{2};\ldots;e_{T};\underbrace{e_{\rm pad};e_{\rm pad};\ldots;e_{\rm pad}}_{T_{\max}-T}]\in\mathbb{R}^{T_{\max}\cdot D}\,,$
@@
Kernel choice. In this work, we focus on the linear and the Gaussian kernel. The linear kernel (Dot in the following) $k^{\text{lin}}(a,b)=a^{\top}b$ is the simplest choice, and gives rise to simple measures of similarity. Indeed, it is easy to show that the MMD in that case corresponds to the $L^{2}$ distance between $y_{\text{beam}}$ and the barycenter of $Y^{\tau}_{\rm sto}$. On the other side, the Gaussian kernel is often preferred in machine learning applications, notably because of the flexibility of its parameter tuning. For any $a,b\in\mathbb{R}^{D}$, it is defined as
@@
In the previous display, the bandwidth parameter $\gamma>0$ controls the scaling of input distances before taking the exponential. The median heuristic (Garreau2018-largesampleanalysismedian, ), which sets the length scale to the median of pairwise distances, is commonly used in kernel methods due to its empirical effectiveness. We explored various settings of length scale selections, and selected the best percentile at $25$-th percentile, which is reported in Appendix F.
To conclude this section, we note that MMD is a good choice when the data dimensionality is high and the available sample size is limited (Gretton-2012a, ). These MMD’s advantages are well-suited to our setting since the vectorized outputs can have quite a high dimension (especially in the concat setting), while the number of sampled sequences $H^{\tau}_{\text{sto}}$ is strongly limited by our compute budget.
@@ 2.3 MMD Flagger
Now, we are equipped with all the machinery required to introduce our proposed method. MMD-Flagger is a simple rule-based approach checking MMD distances $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau}_{\rm sto})$ for a grid of temperature parameters $T=\{\tau_{1},\dots,\tau\}$. By default, we propose to take $T=\{0.1,\ldots,1.0\}$ with a step of $0.1$. For each $\tau\in T$, we query the model to sample token sequences $Y^{\tau}_{\rm sto}:=\{{y^{\tau}_{\rm sto}}\}_{n=1}^{N}$ and its vector representations $H^{\tau}_{\rm sto}:=\{h_{\rm sto}^{\tau}\}_{n=1}^{N}$ by the stochastic sampling with temperature parameter $\tau$, as described in Section 2.1. We then compute the MMD distances defined by Eq. (3). Namely,
$\mathcal{T}_{\widehat{MMD}}:=(\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau=0.1}_{\rm sto}),\ldots,\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau}_{\rm sto}))\,.$
Since an empirical MMD estimate (Eq 3) converges to the true MMD as the sample size increases, the variability in the computed MMD trajectory gradually disappears, as empirically shown in Appendix H.
After computing these values, we detect the inflection point of this trajectory and flag it as a hallucination if it is close to 00. To detect this inflection point, we proceed in the following way: first, we consider a minimal temperature $\tau_{0}$, set by default at $0.11$. Then we compute $\tau_{\text{min}}$ the temperature parameter realizing the minimal MMD distance between the reference and the stochastic samples. In other words, $\tau_{\min}$ is the index of $\mathcal{T}_{\widehat{MMD}}$ with the minimal value. Finally, we flag as hallucination if $\tau_{\min}$ is greater than or equal to $\tau_{0}$. Figure 1 showcases an example of the method at work in both the hallucination and the non-hallucination scenario. We collect the operating procedure in Algorithm 1.
Input: LLM, source input $x$, temperature parameters $T$, number of samples $N$, a minimum temperature for detecting the hallucination $\tau_{0}$
Obtain beam-search translation $h_{\text{beam}}=\text{LLM}_{\text{beam}}(x)$
@@
       Repeat stochastic translation in $N$ times and obtain $H_{\text{sto}}^{\tau}=\text{LLM}_{\text{sto}}^{\tau}(x)$
       Compute the MMD distance $\widehat{\rm MMD}^{2}_{n}(H_{\rm beam},H^{\tau}_{\rm sto})$ using Eq. (3)
@@ 3 Empirical Assessment
In this section, we evaluate MMD-Flagger on two machine translation datasets. All experiments were done on a single NVIDIA Tesla V100 GPU. We report more details on the experimental setup in Appendix B.
Table 2: Hallucination detection performance reported as “Recall (Precision).” We mark the best recall in bold. MMD-Flagger matches existing baseline in terms of recall on the LFAN-HALL dataset, and outperforms in terms of recall and precision on the Halomi dataset. We refer to Section 3.3 for more details.
LFAN-HALL Halomi
TNG 0.069 (1.000) 0.002 (1.000)
Baselines MC-Dsim 0.815 (0.182) 0.553 (0.429)
Seq-Logprob 0.850 (0.189) 0.381 (0.273)
MMD-Flagger (avg-Gaussian) 0.816 (0.129) 0.825 (0.281)
Ours MMD-Flagger (avg-Dot) 0.803 (0.124) 0.779 (0.278)
MMD-Flagger (concat-Gaussian) 0.850 (0.134) 0.824 (0.278)
MMD-Flagger (concat-Dot) 0.833 (0.130) 0.817 (0.282)
@@ 3.1 Datasets and metrics
LFAN-HALL (Guerreiro-2023-looking, ) is a hallucination detection benchmark dataset for the machine translation task from German to English. The dataset consists in $3,413$ machine translations obtained by a transformer-based encoder-decoder model relying on the fairseq package Ott-2019-fairseq and beam search, together with associated reference translation and human-annotated hallucination labels. Since we observed some inaccurate annotation labels, we undertook a review of the annotated labels and made corrections. The relabeling process is detailed in Appendix A. The dataset defines three types of hallucination labels occupying approximately $8\%$ of all translations, and we unify these three labels as “hallucination.” In this dataset, the beam search width is fixed to $5$.
Halomi (Dale-2022-detecting, ) is a hallucination detection benchmark dataset consisting over $3,000$ sentences in $9$ languages. Similarly to LFAN-HALL, Halomi consists of beam search translations manually annotated for hallucination. The dataset defines three types of hallucinations that occupy approximately $26\%$ of the dataset, and we again unify them as “hallucination.” Following the setting of (Himmi-2024-enhanced, ), we use a distilled version of the multilingual translation model (Costa-2022-no, ), totaling 600M parameters. We use the open source Hugging Face implementation. Beam search width was also set to $5$.
@@ 3.2 MMD-Flagger settings
As temperature parameters, we set $T=\{0.1,0.2,\ldots,0.9,1.0\}$. Previous work (Peeperkorn-etal-2024, ) used a smaller value $\tau<0.1$, however, in our setting, the minimum $\tau$ is set to $\tau=0.1$ since the fairseq package runs into numerical problems whenever $\tau<0.1$. The set of stochastic sampling sequences $Y_{\text{sto}}^{\tau}$ and corresponding vector representations $H_{\rm sto}^{\tau}$ consist of $25$ examples for each temperature parameter. We obtain the embedding representation from the word embedding layer of the given LLM. The embedding dimensions are $512$ for LFAN-HALL and $1,024$ for Halomi.
As described in Section 2.2, we considered the linear(Dot) and Gaussian (Gaussian) kernels in our experiments. To pick the bandwidth $\gamma$ for the Gaussian kernel, we randomly select $200$ (correct) translations as calibration data. The $T_{\rm max}$ of concat is set to the maximum token length among $200$ translations. We conduct experiments with various configurations of choosing the length scale and select the $25$-th percentile of $L^{2}$-distance, instead of the usual median heuristic (Garreau2018-largesampleanalysismedian, ). Supporting experiments for this choice can be found in Appendix F. For the Halomi dataset, selecting the calibration data and calculating the length scale is done per source and target language pair.
We also benchmarked MMD-Flagger on internal representations of the model (namely, taking as $h$ vectors of the intermediate hidden layers). We did not find a consistent pattern in our results, while some internal representations yield better experimental results, see Appendix G.
@@ 3.3 Baseline Methods
We considered the baselines proposed by (Guerreiro-2023-looking, ); one rule-based method (TNG) and two uncertainty estimation methods (MC-DSim, Seq-Logprob). Let us describe them briefly.
Top n-gram count (TNG(Raunak-etal-2021-curious, ) is a rule-based method that compares frequencies of $n$-grams between the source and the translation text. TNG flags the translation as the hallucination if certain $n$-gram counts have significant differences. TNG is designed to detect the oscillation type of hallucination; thus, it struggles to flag strong or fully-detached hallucination cases. We set $n=4$ following (Guerreiro-2023-looking, ).
Dissimilarity of Monte-Carlo hypotheses (MC-DSim(Fomicheva-2020-unsupervised, ; Zerva-2021-unbabel, ) is the average similarity scores between $y_{\text{beam}}$ and $Y_{\rm MC}$, defined by $\frac{1}{N}\sum_{i=1}^{N}\text{SIM}(y_{\text{beam}},Y_{\rm MC}),$ where $Y_{\rm MC}$ is a token sequence generated by the beam search using Monte-Carlo Dropout (Gal-2016-dropout, ), and $N$ is the number of Monte-Carlo Dropout generations. Following (Guerreiro-2023-looking, ), we set $N=10$ and use the METEOR metric (Banerjee-2005-meteor, ) of the nltk package (Bird-loper-2004-nltk, ) to compute the similarity score.
Sequence log-probability (Seq-Logprob) is the average of log-probability, defined by $\frac{1}{T}\sum_{k=1}^{T}\log\mathbb{P}\left(y_{k}\middle|y_{<k},x,\theta\right),$ where $x$ is the source text, $y_{k}$ is the $k$-th token in the translation using the beam search, $y_{<k}$ is the previous tokens, $T$ is the length of tokens in the translation, and $\theta$ is the model parameters of a LLM.
MC-DSim and Seq-Logprob are both using a fixed threshold to detect hallucination. As Guerreiro-2023-looking suggests, we fix the threshold at the $40$-th percentile among a given dataset. On the LFAN-HALL dataset, the threshold selection is performed using the whole dataset, while we compute the threshold per language pair of the source and target language on the Halomi dataset. The threshold selection is described in more depth in Appendix C.2.

### v3

@@ 2.2 Maximum Mean Discrepancy (MMD)
Maximum Mean Discrepancy (MMD) is a measure of distance between two probability distributions. It relies on a positive semi-definite kernel (Aronszajn, 1950; Hofmann et al., 2008; Smola et al., 2009) to map the distributions into a new space, a Reproducing Kernel Hilbert Space (RKHS), where probability distributions are represented as data points (Muandet et al., 2017). A large MMD value indicates significant differences between two distributions. We depict an intuitive presentation of the MMD in Figure 2.
@@
[…]
@@
Empirical estimate. Given i.i.d. samples ${\bf A}=\{A^{1},\dots,A^{n}\}\sim P$ and ${\bf B}=\{B^{1},\dots,B^{m}\}\sim Q$, one can estimate the MMD by the corresponding empirical averages:
$\begin{split}\widehat{\rm MMD}^{2}_{U}({\bf A},{\bf B})&:=\frac{1}{n(n-1)}\sum_{\begin{subarray}{c}1\leq i,i^{\prime}\leq n\\ i\neq i^{\prime}\end{subarray}}k(A^{i},A^{i^{\prime}})\\ +\frac{1}{m(m-1)}\sum_{\begin{subarray}{c}1\leq j,j^{\prime}\leq m\\ j\neq j^{\prime}\end{subarray}}&k(B^{j},B^{j^{\prime}})-\frac{2}{nm}\sum_{\begin{subarray}{c}i=1,\dots,n\\ j=1,\dots,m\end{subarray}}k(A^{i},B^{j}).\end{split}$ (1)
This is an unbiased estimate of the MMD between two distributions $P$ and $Q$ and converges at the rate $\mathcal{O}_{\mathbf{P}}(\min(n,m)^{-1/2})$ as $n,m\to\infty$ provided that $\sup_{x\in\mathbb{R}^{d}}k(x,x)<\infty$ (Gretton et al., 2012, Theorem 10).
MMD in our setting. Recall that our goal is to measure the similarity between the outputs of $\text{LLM}_{\rm hyp}(x)$ and those of $\text{LLM}_{\rm sto}^{\tau}(x)$. To do this, we consider these two as two probability distributions, and use Eq. (1) to get an estimate of the MMD.
Vector representation of outputs. To capture the rich semantic relationships between generated outputs, we convert the sequences into dense token embedding vectors. The simplest representation is to resort to the word embedding matrix $E\in\mathbb{R}^{V\times D}$ used by LLMs. Thus, each token sequence $y$ is associated with a sequence of embedding vectors
@@
To make it a fixed-length vector $h$, we introduce two aggregations of $T$-length vector representations; avg and concat. The avg aggregation is computing the average over $T$ vectors, namely
@@
$h=[e_{1};e_{2};\ldots;e_{T};\underbrace{e_{\rm pad};\ldots;e_{\rm pad}}_{T_{\max}-T}]\in\mathbb{R}^{T_{\max}\cdot D}\,$,
@@
Kernel choice. We briefly introduce the linear and Gaussian kernel. Choices of other kernels are discussed and evaluated in Appendix D.2. The linear kernel (Dot in the following) $k^{\text{lin}}(a,b)=a^{\top}b$ is the simplest choice, and gives rise to simple measures of similarity. Indeed, it is easy to show that the MMD in that case corresponds to the $L^{2}$ distance between $y_{\text{beam}}$ and the barycenter of $Y^{\tau}_{\rm sto}$. On the other side, the Gaussian kernel is often preferred in machine learning applications, notably because of the flexibility of its parameter tuning. For any $a,b\in\mathbb{R}^{D}$, it is defined as
@@
In the previous display, the bandwidth parameter $\gamma>0$ controls the scaling of input distances before taking the exponential. The median heuristic (Garreau et al., 2017), which sets the length scale to the median of pairwise distances, is commonly used in kernel methods due to its empirical effectiveness. We explored various settings of length scale selections, and selected the best percentile at $25$-th percentile, which is reported in Appendix D.1.
To conclude this section, we note that MMD is a good choice when the data dimensionality is high and the available sample size is limited (Gretton et al., 2012). These MMD’s advantages are well-suited to our setting since the vectorized outputs can have quite a high dimension (especially in the concat setting), while the number of sampled sequences $H^{\tau}_{\text{sto}}$ is strongly limited by our compute budget.
@@ 2.3 MMD Flagger
Now, we are equipped with all the machinery required to introduce our proposed method. MMD-Flagger is a simple rule-based approach checking MMD distances $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau}_{\rm sto})$ for a grid of temperature parameters $T=\{\tau_{1},\dots,\tau\}$. By default, we propose to take $T=\{0.1,\ldots,1.0\}$ with a step of $0.1$. For each $\tau\in T$, we query the model to sample token sequences $Y^{\tau}_{\rm sto}:=\{{y^{\tau}_{\rm sto}}\}_{n=1}^{N}$ and its vector representations $H^{\tau}_{\rm sto}:=\{h_{\rm sto}^{\tau}\}_{n=1}^{N}$ by the stochastic sampling with temperature parameter $\tau$, as described in Section 2.1. We then compute the MMD distances defined by Eq. (1). Namely,
$\begin{split}\mathcal{T}_{\widehat{MMD}}:=(&\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau=0.1}_{\rm sto}),\\ &\ldots,\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau}_{\rm sto}))\,.\end{split}$
Since an empirical MMD estimate (Eq 1) converges to the true MMD as the sample size increases, the variability in the computed MMD trajectory gradually disappears, as empirically shown in Appendix F.
After computing these values, we detect the inflection point of this trajectory and flag it as a hallucination if it is close to 0. To detect this inflection point, we proceed in the following way: first, we consider a minimal temperature $\tau_{0}$, set by default at $0.11$. Then we compute $\tau_{\text{min}}$ the temperature parameter realizing the minimal MMD distance between the reference and the stochastic samples. In other words, $\tau_{\min}$ is the index of $\mathcal{T}_{\widehat{MMD}}$ with the minimal value. Finally, we flag as hallucination if $\tau_{\min}$ is greater than or equal to $\tau_{0}$. Figure 1 showcases an example of the method at work in both the hallucination and the non-hallucination scenario. We collect the operating procedure in Algorithm 1.
Input: LLM, source input $x$, temperature parameters $T$, number of samples NN, a minimum temperature for detecting the hallucination $\tau_{0}$
Obtain a hypothesis sequence $H_{\text{hyp}}$
@@
   Repeat stochastic translation in NN times and obtain $H_{\text{sto}}^{\tau}=\text{LLM}_{\text{sto}}^{\tau}(x)$
   Compute the MMD distance $\widehat{\rm MMD}^{2}_{n}(H_{\rm hyp},H^{\tau}_{\rm sto})$ using Eq. (1)
@@ 3 Empirical Assessment
In this section, we evaluate MMD-Flagger on four hallucination benchmark datasets. All experiments were done on a single NVIDIA Tesla V100 GPU. We report more details on the experimental setup in Appendix A and comparison with SelfCheckGPT (Manakul et al., 2023) leveraging external resources in Appendix C.1.
Table 2: Hallucination detection performance reported as “Recall (Precision).” We mark the best recall in bold. LFAN-HALL, Halomi are machine translation datasets, and ConstFact, FaithBench are summarization datasets.
LFAN-HALL Halomi ConstFact FaithBench
TNG 0.069 (1.000) 0.002 (1.000) - -
Baselines MC-Dsim 0.815 (0.182) 0.553 (0.429) 0.458 (0.750) 0.363 (0.266)
Seq-Logprob 0.850 (0.189) 0.381 (0.273) 0.348 (0.570) 0.363 (0.4)
avg-Gaussian 0.816 (0.129) 0.825 (0.281) 0.904 (0.667) 0.909 (0.555)
Ours avg-Dot 0.803 (0.124) 0.779 (0.278) 0.872 (0.677) 0.909 (0.625)
concat-Gaussian 0.850 (0.134) 0.824 (0.278) 0.879 (0.648) -
concat-Dot 0.833 (0.130) 0.817 (0.282) 0.838 (0.644) -
@@ 3.1 Datasets and metrics
We evaluate MMD-Flagger and baselines on four distinct hallucination benchmarks; two for the machine translation task LFAN-HALL (Guerreiro et al., 2023b), Halomi (Dale et al., 2023) and the other two for the summarization task ConstFact (Dreyer et al., 2023), FaithBench (Bao et al., 2025).
LFAN-HALL dataset is a machine translation hallucination detection benchmark. The dataset was generated using an encoder-decoder model provided by the authors, that is a smaller version of a WMT18 translation model with roughly 7777M parameters, produced the $3,413$ translations in the dataset. We note that after our manual review of the annotations, the corrected dataset, included in the supplementary materials, reveals that approximately $8\%$ of the translations are classified as hallucinated.
Halomi dataset is a multilingual hallucination detection benchmark for machine translation, encompassing over $3,000$ sentences across nine languages. The translations were generated using the multilingual encoder-decoder model nllb-200-distilled-600M (Costa-Jussà et al., 2022). In this dataset, approximately $26\%$ of the translations are annotated as hallucinated.
ConstFact dataset evaluates the factual consistency of summaries derived from the CNN/Daily Mail dataset. The summaries within ConstFact were generated using the BART encoder-decoder architecture (Lewis et al., 2020), specifically the bart.large.cnn variant with over 406406M parameters.
FaithBench dataset is a benchmark of evaluating the faithfulness of summaries from decoder-only models. The hypothesis outputs are generated with argmax strategy that does not fit with MMD-Flagger, discussed in Section 4. To tailor this dataset to our needs, we constructed a custom sub-dataset. We randomly selected $25$ source documents and regenerated summaries using the beam search decoding strategy with the Phi-3-mini-4k model (Grattafiori et al., 2024). This is in contrast to the original dataset’s use of the argmax decoding. The model has 3.83.8B parameters. Our newly annotated sub-dataset, with $11$ hallucinated and $14$ faithful summaries, is included in our supplementary materials.
@@ 3.2 MMD-Flagger settings
As temperature parameters, we set $T=\{0.1,0.2,\ldots,0.9,1.0\}$. The set of stochastic sampling sequences $Y_{\text{sto}}^{\tau}$ and corresponding vector representations $H_{\rm sto}^{\tau}$ consist of $25$ examples for each temperature parameter. We obtain the embedding representation from the word embedding layer of the given LLM. The embedding dimensions are $512$ for LFAN-HALL and $1,024$ for Halomi and ConstFact, and $3,072$ for FaithBench. We do not evaluate a vector representation concat of MMD-Flagger on FaithBench due to the computational constraints that the concatenated vector becomes too large.
As described in Section 2.2, we considered the linear (Dot) and Gaussian (Gaussian) kernels in our experiments. To pick the bandwidth $\gamma$ for the Gaussian kernel, we randomly select $200$ (correct) translations as calibration data. The $T_{\rm max}$ of concat is set to the maximum token length among $200$ translations. We conduct experiments with various configurations of choosing the length scale and select the $25$-th percentile of $L^{2}$-distance, instead of the usual median heuristic (Garreau et al., 2017). Supporting experiments for this choice can be found in Appendix D.1. For the Halomi dataset, selecting the calibration data and calculating the length scale is done per source and target language pair.
We also benchmarked MMD-Flagger on internal representations of the model (namely, taking as $h$ vectors of the intermediate hidden layers). We did not find a consistent pattern in our results, while some internal representations occasionally yield better experimental results, see Appendix E.
@@ 3.3 Baseline Methods
We considered the baselines proposed by (Guerreiro et al., 2023b); one rule-based method (TNG) and two uncertainty estimation methods (MC-DSim, Seq-Logprob). Let us describe them briefly.
Top n-gram count (TNG(Raunak et al., 2021) is a rule-based method that compares frequencies of $n$-grams between the source and the translation text. TNG flags the translation as the hallucination if certain $n$-gram counts have significant differences. TNG is designed to detect the oscillation type of hallucination; thus, it struggles to flag strong or fully-detached hallucination cases. We set $n=4$ following (Guerreiro et al., 2023b). The evaluation of TNG is omitted on ConstFact and FaithBench datasets as it targets repetition-type hallucinations not present in these datasets.
Dissimilarity of Monte-Carlo hypotheses (MC-DSim(Fomicheva et al., 2020; Zerva et al., 2021) is the average similarity scores between $y_{\text{hyp}}$ and $Y_{\rm MC}$, defined by $\frac{1}{N}\sum_{i=1}^{N}\text{SIM}(y_{\text{hyp}},Y_{\rm MC}),$ where $Y_{\rm MC}$ is a token sequence generated by the beam search using Monte-Carlo Dropout (Gal and Ghahramani, 2016), and $N$ is the number of Monte-Carlo Dropout generations. Following (Guerreiro et al., 2023b), we set $N=10$ and use the METEOR metric (Banerjee and Lavie, 2005) of the nltk package (Bird and Loper, 2004) to compute the similarity score.
Sequence log-probability (Seq-Logprob) is the average of log-probability, defined by $\frac{1}{T}\sum_{k=1}^{T}\log\mathbb{P}\left(y_{k}\middle|y_{<k},x,\theta\right),$ where $x$ is the source text, $y_{k}$ is the $k$-th token in the translation using the beam search, $y_{<k}$ is the previous tokens, $T$ is the length of tokens in the translation, and $\theta$ is the model parameters of the LLM.
MC-DSim and Seq-Logprob are both using a fixed threshold to detect hallucination. As Guerreiro et al. (2023b) suggests, we fix the threshold at the $40$-th percentile among a given dataset. On the LFAN-HALL dataset, the threshold selection is performed using the whole dataset, while we compute the threshold per language pair of the source and target language on the Halomi dataset. The threshold selection is described in more depth in Appendix C.